Choose the proportion you would use to solve the problem below.
A truck can pump 800 gallons of oil in 35 minutes. How much oil can it pump in 2 hours?

can u type just how u solved it




Answer by edjones(7311)

x/120=800/35
120x/120=120*800/35
x=96000/35
x=2742.86 gal
.
Ed